OpenStudy (anonymous):

Solve. 3x2 + 4x = −5 the quantity of negative 2 plus or minus i square root 11 all over 3 the quantity of negative 2 plus or minus 4i square root 3 all over 3 the quantity of negative 2 plus or minus 2 square root 11 all over 3 the quantity of negative 2 plus or minus 4i square root 3 all over 3 i will reward medal

OpenStudy (therealmeeeee):

3x^2+4x+5=0

OpenStudy (therealmeeeee):

3x^2 + 4x = –5 3x^2 +4x +5 =0 Now, use the quadratic formula for the equation. x=−b±b2−4ac−−−−−−−√2a Put a =3, b=4 and c=5 into the formula. Can you do it?

OpenStudy (anonymous):

yea i can.thanks

OpenStudy (therealmeeeee):

thanks

OpenStudy (anonymous):

welcome